OpenStudy (anonymous):

Two lines, C and D, are represented by the equations given below: Line C: y = x + 14 Line D: y = 3x + 2 Which of the following shows the solution to the system of equations and explains why?

OpenStudy (anonymous):

BY THE WAY I GIVE MEDALS

OpenStudy (danjs):

If y = x +14 and y = 3x+2 set y = y

OpenStudy (danjs):

x+14 = 3x + 2 solve for x, then use that value of x to find y

OpenStudy (danjs):

Here is the graph of the two lines. The intersection point is the solution x= y= you are looking for (point (x,y)